OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–2, 1) and (4, 13)? please help me

OpenStudy (ivettef365):

what is the slope ?

OpenStudy (anonymous):

y – 1 = –2(x – 4) y – 1 = 2(x + 2) y – 13 = –2(x + 2) y – 13 = 2(x + 4)

OpenStudy (ivettef365):

do you know what is the slope ?

OpenStudy (anonymous):

no I dont unfortunatly

OpenStudy (ivettef365):

ok slope you use formula y2 - y1 ------ x2 - x1 you have points (-2,1) and (4,13) , can you plug in on formula (x1,y1) (x2,y2)

OpenStudy (anonymous):

i still dont get how to get to the answer after pluggin in and solving

OpenStudy (ivettef365):

13 - 1 12 ----- = ---- = 2 4 + 2 6 so the slope = 2

OpenStudy (ivettef365):

the formula for the equation in point slope form is y- y1 = m(x-x1) you have m= slope = 2, use points (-2,1) for (x1,y1) what do you get ?

OpenStudy (ivettef365):

can you plug numbers in this formula, that will give you answer

OpenStudy (anonymous):

y – 1 = 2(x + 2) ?

OpenStudy (ivettef365):

correct